# Fast and Accurate Decoder for the XZZX code Using Simulated Annealing

### v1

@@ IV Decoding algorithms
We elaborate on our SA decoder in Sec. IV.1. Subsequently, we describe the MWPM decoder in Sec. IV.2. As a simplified version, in Sec. IV.3, we propose a greedy matching decoder, exploited to construct an initial configuration of the SA decoder.
@@ IV.1 Simulated annealing (SA) decoder
We assume the code capacity noise where only the data qubits suffer errors. Specifically, each data qubit suffers an error of $X$-, $Y$-, or $Z$-type, independently and identically obeying the probabilities $p_{x}$, $p_{y}$, $p_{z}$, respectively. Any type of error occurs with the probability $p\coloneqq p_{x}+p_{y}+p_{z}$, called physical error probability. This noise for each data qubit $\rho$ is expressed as quantum channel
@@
For an error configuration $C$, let $n_{x},n_{y},n_{z}$ denote the numbers of qubits which suffer $X,Y,Z$ errors, respectively. The probability of $C$ is then expressed as
$\operatorname{Prob}(C)$ $\coloneqq$ ${p_{x}}^{n_{x}}\,{p_{y}}^{n_{y}}\,{p_{z}}^{n_{z}}\,(1-p)^{N_{\mathrm{q}}-(n_{x}+n_{y}+n_{z})}$ (6)
@@
$=$ $\left(\frac{p}{1-p}\right)^{\alpha_{x}n_{x}+\alpha_{y}n_{y}+\alpha_{z}n_{z}}$ (8)
is a factor relevant to $p_{x},p_{y},p_{z}$. From Eqs. (7),(8), the coefficients $\alpha_{x},\alpha_{y},\alpha_{z}$ are expressed as
$\alpha_{\mu}=\frac{\log{\frac{p_{\mu}}{1-p}}}{\log{\frac{p}{1-p}}}$ (9)
where $\mu\in\{x,y,z\}$. Assuming that we know or estimate $p_{x}$, $p_{y}$, $p_{z}$ enables optimal decoding, as discussed below.
We define energy of the error configuration $C$ as the weighted sum of the numbers [^2] The numbers $n_{x}$, $n_{y}$, $n_{z}$ are calculated as follows. We formulate and implement the Metropolis method with bit representation taking values in the set $\{0,1\}$, instead of Ising spins taking values in the set $\{1,-1\}$. In an error configuration, an error that suffers each data qubit is represented by two bits. In the two bits, the upper and lower bit means for $X$ and $Z$ error, respectively. Thus, $I$, $Z$, $X$, $Y$ errors are represented as $00$, $01$, $10$, $11$, respectively. As such, counting the numbers of qubits which have bit representations $01$, $10$, $11$ yields $n_{z}$, $n_{x}$, $n_{y}$, respectively. :
@@
$\pi(C)=\left(\frac{p}{1-p}\right)^{\mathcal{H}(C)}=e^{-\beta\,\mathcal{H}(C)}$ (11)
where $\beta$ is inverse temperature, defined as [^3] The Nishimori temperature includes the coefficient $1/2$ in the formulation by Ising spins [28, 29]. Instead, we employ the formulation by bit representation, and hence, this coefficient does not appear.
$\beta\coloneqq-\log\left(\frac{p}{1-p}\right).$ (12)
Let us define an equivalence class for an error configuration $C$ via the equivalence class $\cdot\,$]. Given a syndrome $S$, we can construct a pure error, i.e., an error configuration $T(S)$ consistent with the syndrome $S$. Using any pure error $T(S)$, any error configuration $C$ consistent with the syndrome $S$ can be decomposed as
$C=T(S)\,L_{P}\,G$ (13)
where $P\in\{I,X,Y,Z\}$ and a stabilizer operator $G\in\mathcal{G}$ and vice versa [30].
Left-multiplying both sides of Eq. (13) by $T(S)$ yields
$T(S)\,C]=[L_{P}].$ (14)
Then, the set of error configurations $C$ satisfying Eq. (14) is defined as
$\operatorname{Rep}_{T(S)}(L_{P})$ $\coloneqq$ $\left\{C\,\middle|\,[T(S)\,C]=[L_{P}]\right\}$ (15)
$=$ $\left\{T(S)\,L_{P}\,G\,\middle|\,G\in\mathcal{G}\right\},$ (16)
which is the set of representatives of another equivalence class, referred to as class for the logical operator $L_{P}$ based on $T(S)$. We employ a pure error $T(S)$ as output of some decoder. Specifically, we utilize our greedy matching decoder, elaborated in Sec. IV.3.
The probability of the class for the logical operator $L_{P}$ based on $T(S)$ is given by
$\mathrm{P}_{L_{P}}\coloneqq\max_{C}\left\{\operatorname{Prob}(C)\,\middle|\,C\in\operatorname{Rep}_{T(S)}(L_{P})\right\}.$ (17)
Eq. (17) is expressed by using Eqs. (6),(11) as
$\mathrm{P}_{L_{P}}\propto e^{-\beta\,E_{L_{P}}}$ (18)
@@
$E_{L_{P}}$ $\coloneqq$ $\min_{C}\left\{\mathcal{H}(C)\,\middle|\,C\in\operatorname{Rep}_{T(S)}(L_{P})\right\}$ (19)
$=$ $\min_{G\in\mathcal{G}}\mathcal{H}\left(T(S)\,L_{P}\,G\right)$
@@
We determine the most likely class for the logical operator $L_{P^{*}}$ which maximizes the error probabilities $\mathrm{P}_{L_{P}}$:
$P^{*}\coloneqq\operatorname*{\arg\max}_{P\in\{I,X,Y,Z\}}\mathrm{P}_{L_{P}}=\operatorname*{\arg\min}_{P\in\{I,X,Y,Z\}}E_{L_{P}},$ (20)
@@
which follows from Eq. (18). This decoding principle is based on the theory of the domain-wall of spin-glass models [31, §3.5] [32, 33]. For each $P\in\{I,X,Y,Z\}$, to calculate $E_{L_{P}}$ in Eq. (19), we employ simulated annealing (SA) [21].
For our SA, in the decomposition of Eq. (13), $C_{L_{P}}\coloneqq T(S)\,L_{P}$ is set as an initial configuration of SA, and the stabilizer operator $G$ is updated by the Metropolis method to explore the ground state energy $E_{L_{P}}$. Although it is rigorously proved that SA converges to the optimal solution regardless of the initial configuration if we take an infinite time [34], we expect that in realistic finite time, the initial configuration affects the convergence speed of SA. We examine this effect in Sec. V.3.
@@
We give our SA decoder in Algorithm 1. To acquire the lowest possible energy $E_{L_{P}}$, we carry out simulated annealing multiple times (say $N_{\mathrm{SA}}$ times) for the identical error configuration $C_{L_{P}}$ in Lines 78 of Algorithm 1, and select the minimum energy among them in Line 9. In Line 14, we finally obtain the recovery configuration $C_{L_{P^{*}}}=T\,L_{P^{*}}$. [^4] As a recovery error configuration, any configuration $C\in\operatorname{Rep}_{T(S)}(L_{P^{*}})$ can be selected.
@@
It should be noted that, in the nested for loops in Lines 5 and 7 in Algorithm 1, each iteration is an independent SA task and, hence, these tasks are obviously parallelizable. The total number of iterations in these loops are $4$ and $N_{\mathrm{SA}}$ and, therefore, if parallelized, the elapsed time in parallel execution is expected to be the elapsed time in sequential execution divided by $4N_{\mathrm{SA}}$. This parallelized implementation is left to our future work.
Algorithm 1 invokes the simulated annealing, shown in Algorithm 2. Algorithm 2 performs the Metropolis method, given in Algorithm 3. Line 9 in Algorithm 3 flips spins, corresponding to errors on qubits, by acting a stabilizer generator [^5] These spins are implemented by bits. Hence, the multiplication of error configurations $C\,G_{f}$ is implemented as addition modulo two. in order to keep satisfying the syndrome consistency condition $\partial{C}=S$ throughout the entire SA algorithm.
For the physical error probability $p$, Eq. (12) gives the Nishimori temperature $\beta_{\mathrm{nishimori}}$ where actual error configuration and recovery configuration equally fluctuate, and we can incorporate the degeneracy of error configurations with the same energy. Hence, it is optimal for decoding as a fixed temperature [28, 29, 33].
@@
In Algorithm 2, to accelerate convergence of SA, we employ inverse temperature scheduling with logarithmic function from $\beta_{\mathrm{init}}\coloneqq 0.9\,\beta_{\mathrm{nishimori}}$ to $\beta_{\mathrm{final}}\coloneqq\beta_{\mathrm{nishimori}}$
@@
$\beta_{i}\coloneqq\beta_{\mathrm{init}}\,\left(1+r\,\log{i}\right)\quad\text{for $i=1,\dots,N_{\mathrm{\beta}}$}$ (21)
@@
$r\coloneqq\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\left(\log{N_{\mathrm{\beta}}}\right)\right.$. (22)
For each inverse temperature $\beta$, we execute the Metropolis step $N_{\mathrm{stabilizer}}$ times, which is counted as one Monte Carlo step, according to the convention.
@@
1: $S$ : Syndrome for error configuration
2: Recovery configuration $C$
@@
4:   $T\leftarrow$ greedy_matching($S$) $\triangleright$ pure error
5:   for all $P\in\{I,X,Z,Y\}$ do
6:    $C_{L_{P}}\coloneqq T\,L_{P}$ $\triangleright$ Initial configuration for SA
@@
8:      $E^{(i)}_{L_{P}}\leftarrow$ simulated_annealing($C_{L_{P}}$)
9:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}}$
@@
12:   $P^{*}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (20) 13:
14:   return $C_{L_{P^{*}}}$
@@
2: Ground state energy (or lower excited state energy close to the ground state energy)
@@
5:   $E_{\mathrm{min}}\leftarrow E$
6:   $r\coloneqq\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\left(\log{N_{\mathrm{\beta}}}\right)\right$. $\triangleright$ Eq. (22) 7:
8:   for $i=1$ to $N_{\mathrm{\beta}}$ do
9:    $\beta\leftarrow\beta_{\mathrm{init}}\,\left(1+r\,\log{i}\right)$ $\triangleright$ Eq. (21)
10:    for $N_{\mathrm{stabilizer}}$ times do
11:      $\Delta E\leftarrow$ metropolis_Step($\beta$,  $C$)
12:      $E\leftarrow E+\Delta E$
13:      $E_{\mathrm{min}}\leftarrow\min\{E_{\mathrm{min}},\,E\}$ $\triangleright$ Update min energy
14:
15:   return $E_{\mathrm{min}}$
Algorithm 3 Metropolis method which acts a stabilizer generator
1: $\beta$ : inverse temperature,  $C$ : Error configuration
2: Difference of energy of error configurations
3: function Metropolis_Step($\beta$,  $C$)
4:   Select a stabilizer generator $G_{f}$ randomly.
5:   $\Delta E\leftarrow\mathcal{H}(C\,G_{f})-\mathcal{H}(C)$ 6:
7:   Generate a uniform random probability $r\in[0,1]$.
8:   if $\Delta E<0$ or $r<e^{-\beta\,\Delta E}$ then
9:    $C\leftarrow C\,G_{f}$ $\triangleright$ Flips spins
10:    return $\Delta E$
11:   else
12:    return 0

### v2

@@ IV Decoding algorithms
We elaborate on our SA decoder in Sec. IV.1. Subsequently, we describe the MWPM decoder in Sec. IV.2. As a simplified version, in Sec. IV.3, we propose a greedy matching decoder, exploited to construct an initial configuration for the SA decoder.
@@ IV.1 Simulated annealing (SA) decoder
We assume the code capacity noise model where only data qubits are noisy and syndrome measurements are perfect. Specifically, each data qubit suffers an error of $X$-, $Y$-, or $Z$-type, independently and identically distributed with the probabilities $p_{x}$, $p_{y}$, $p_{z}$, respectively. A non-identity Pauli error occurs with the probability $p\coloneqq p_{x}+p_{y}+p_{z}$, called the physical error rate. The noise acting on each data qubit state $\rho$ is given by the channel
@@
For an error configuration $C$, let $n_{x},n_{y},n_{z}$ denote the numbers of qubits which suffer $X,Y,Z$ errors, respectively. The probability of $C$ is then given by
$\Pr(C)$ $\coloneqq$ ${p_{x}}^{n_{x}}\,{p_{y}}^{n_{y}}\,{p_{z}}^{n_{z}}\,(1-p)^{N_{\mathrm{q}}-(n_{x}+n_{y}+n_{z})}$ (6)
@@
is the (unnormalized) Boltzmann weight of configuration $C$, i.e., $\Pr(C)$ up to the constant factor $(1-p)^{N_{\mathrm{q}}}$.
For convenience, we re-express Eq. (7) using a common base p/(1p)p/(1-p):
$\pi(C)$ $=$ $\left(\frac{p}{1-p}\right)^{\alpha_{x}n_{x}+\alpha_{y}n_{y}+\alpha_{z}n_{z}}$ (8)
with
$\alpha_{\mu}\coloneqq\frac{\log\!{\left(\frac{p_{\mu}}{1-p}\right)}}{\log\!{\left(\frac{p}{1-p}\right)}}$ (9)
where $\mu\in\{x,y,z\}$. We assume $0<p<1/2$, so that $\log(p/(1-p))<0$. Knowing or accurately estimating $p_{x}$, $p_{y}$, $p_{z}$ enables optimal minimum-energy (MAP-configuration) decoding under the dominant-term approximation discussed below.
We define the energy of the error configuration $C$ as the weighted sum of the numbers [^1] The numbers $n_{x}$, $n_{y}$, $n_{z}$ are calculated as follows. We formulate and implement the Metropolis method with bit representation taking values in the set $\{0,1\}$, instead of Ising spins taking values in the set $\{1,-1\}$. In an error configuration, the error on each data qubit is represented by two bits. The upper and lower bits represent an $X$ and a $Z$ error, respectively. Thus, $I$, $Z$, $X$, $Y$ errors are represented as $00$, $01$, $10$, $11$, respectively. As such, counting the numbers of qubits which have bit representations $01$, $10$, $11$ yields $n_{z}$, $n_{x}$, $n_{y}$, respectively. :
@@
$\pi(C)=\left(\frac{p}{1-p}\right)^{\mathcal{H}(C)}=e^{-\beta_{\mathrm{N}}\,\mathcal{H}(C)},$ (11)
@@
$\beta_{\mathrm{N}}\coloneqq-\log\!\left(\frac{p}{1-p}\right)\qquad\text{for }0<p<1/2.$ (12)
We refer to $\beta_{\mathrm{N}}$ as the Nishimori inverse temperature corresponding to the physical error rate $p$. Conditioned on the syndrome-consistency $\partial C=S$, we consider the posterior $\Pr(C\mid S)\propto\pi(C)$, which follows from Eq. (6).
Given a syndrome $S$, let
$\Omega_{S}\coloneqq\left\{C\in\overline{\mathcal{P}}\,\middle|\,\partial C=S\right\}$ (13)
denote the set of syndrome-consistent configurations. Fix $R(S)\in\Omega_{S}$, referred to as a reference configuration [^2] Such a fixed $R(S)\in\Omega_{S}$ is often called a pure error in the literature [fujii_book].. Then, any error configuration $C\in\Omega_{S}$ can be decomposed as
$C=R(S)\,L_{P}\,G$ (14)
for some $P\in\{I,X,Y,Z\}$ and some $G\in\mathcal{G}$ [Poulin].
Left-multiplying both sides of Eq. (14) by $R(S)$ and using $R(S)^{2}=I$ (up to phase) gives
$R(S)\,C=L_{P}\,G.$ (15)
The product R(S)CR(S)\,C has a trivial syndrome (i.e., $\partial(R(S)\,C)=\emptyset$), so its equivalence class [R(S)C][R(S)\,C] is well-defined (as in Sec. II) and equals one of $L_{P}]$ with $P\in\{I,X,Y,Z\}$. We then define, for each $P\in\{I,X,Y,Z\}$, a coset
$\operatorname{\mathcal{C}}_{R(S)}(L_{P})\coloneqq\left\{C\in\Omega_{S}\,\middle|\,[R(S)\,C]=[L_{P}]\right\}.$ (16)
Note that $L_{P}]$ is an equivalence class in $\overline{\mathcal{P}}/\mathcal{G}$, whereas $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ is a set of concrete configurations $C\in\Omega_{S}$; these objects live in different spaces. When $S=\emptyset$ and we choose $R(S)=I^{\otimes N_{\mathrm{q}}}$, for each $P\in\{I,X,Y,Z\}$, $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ corresponds to the logical operator class $L_{P}]$.
The four cosets $\{\operatorname{\mathcal{C}}_{R(S)}(L_{P})\}_{P\in\{I,X,Y,Z\}}$ form a partition of $\Omega_{S}$:
$\Omega_{S}=\operatorname{\mathcal{C}}_{R(S)}(L_{I})\,\dot{\cup}\,\operatorname{\mathcal{C}}_{R(S)}(L_{X})\,\dot{\cup}\,\operatorname{\mathcal{C}}_{R(S)}(L_{Y})\,\dot{\cup}\,\operatorname{\mathcal{C}}_{R(S)}(L_{Z}),$
[…]
$\operatorname{\mathcal{C}}_{R(S)}(L_{P})=\left\{R(S)\,L_{P}\,G\,\middle|\,G\in\mathcal{G}\right\}.$ (17)
We define the maximum (unnormalized) Boltzmann weight within each coset as
$\pi^{\max}_{L_{P}}\coloneqq\max_{C\in\operatorname{\mathcal{C}}_{R(S)}(L_{P})}\pi(C).$ (18)
A configuration $C$ that attains the maximum in Eq. (18) is the most likely configuration in the coset $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ under the posterior conditioned on $\partial C=S$. Using Eq. (11), Eq. (18) is expressed as
$\pi^{\max}_{L_{P}}=e^{-\beta_{\mathrm{N}}\,E^{\min}_{L_{P}}}$ (19)
where
$E^{\min}_{L_{P}}$ $\coloneqq$ $\min_{C\in\operatorname{\mathcal{C}}_{R(S)}(L_{P})}\mathcal{H}(C)$ (20)
$=$ $\min_{G\in\mathcal{G}}\mathcal{H}\left(R(S)\,L_{P}\,G\right)$
@@
We determine $P^{\ast}$ by maximizing $\pi^{\max}_{L_{P}}$ (equivalently, minimizing $E^{\min}_{L_{P}}$):
$P^{\ast}\coloneqq\operatorname*{\arg\max}_{P\in\{I,X,Y,Z\}}\pi^{\max}_{L_{P}}=\operatorname*{\arg\min}_{P\in\{I,X,Y,Z\}}E^{\min}_{L_{P}},$ (21)
@@
which follows from Eq. (19). This decoding formulation is based on domain-wall constructions in spin-glass models [fujii_book, §3.5] [hukushima_domain_wall, Chubb_stat_model].
Ideally, one would perform maximum-likelihood (ML) decoding over logical operator cosets by comparing the coset weights. We instead use the maximum a posteriori (MAP): we approximate each coset weight $\sum_{C\in\operatorname{\mathcal{C}}_{R(S)}(L_{P})}\pi(C)$ by its largest term $\pi^{\max}_{L_{P}}$ defined in Eq. (19). This reduces ML decoding over cosets to a minimum-energy search within each coset.
For each $P\in\{I,X,Y,Z\}$, to estimate $E^{\min}_{L_{P}}$ in Eq. (20), we employ simulated annealing (SA) [Kirkpatrick]. For our SA, in the decomposition of Eq. (14), $C_{L_{P}}\coloneqq R(S)\,L_{P}$ is set as an initial configuration for SA, and the stabilizer component $G$ is updated by the Metropolis method to search for the ground state energy $E^{\min}_{L_{P}}$.
In most of our simulations, we take $R(S)$ to be the recovery configuration returned by our greedy matching decoder (Sec. IV.3). Although it has been rigorously proved that SA converges to the optimal solution (under suitable cooling schedules) in the infinite-time limit [Geman_Geman], regardless of the initial configuration, in practical finite time, the initial configuration can significantly affect the convergence speed of SA. We examine this effect in Sec. V.3.
@@
We give our SA decoder in Algorithm 1. To estimate $E^{\min}_{L_{P}}$ effectively, we run simulated annealing multiple times (say, $N_{\mathrm{SA}}$ times) starting from the identical error configuration $C_{L_{P}}$ (Lines 78 of Algorithm 1), and select the minimum energy across the runs (Line 9). In Line 14, we output the recovery configuration $C_{L_{P^{\ast}}}=R\,L_{P^{\ast}}$. [^3] We use SA only to estimate $E^{\min}_{L_{P}}$; the output recovery configuration is chosen as $R\,L_{P^{\ast}}$, since any configuration of the inferred logical coset $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ is equivalent up to stabilizers.
@@
Note that, in the nested for loops in Lines 5 and 7 in Algorithm 1, each iteration is an independent SA task and, hence, these tasks are trivially parallelizable. Implementing this parallelization (and quantifying the overhead) is left for future work.
Algorithm 1 invokes the simulated annealing, shown in Algorithm 2. Note that the simulated_annealing function is called with a copy of the initial configuration. Algorithm 2 performs the Metropolis method, given in Algorithm 3.
In Line 10 of Algorithm 3, we update the configuration $C$ by applying a randomly chosen stabilizer generator $G_{f}$.  [^4] We formulate error configurations by bits, instead of Ising spins. Hence, multiplying configurations $C\mapsto C\,G_{f}$ is implemented as bitwise XOR (addition modulo two) on the two-bit representation per qubit. This update keeps the syndrome $S$ invariant.
For simulated annealing (SA), we introduce a tempered family of unnormalized weights
$\pi_{\beta}(C)\coloneqq e^{-\beta\,\mathcal{H}(C)},$ (22)
where $\beta>0$ is treated as a tunable inverse temperature. In particular, the unnormalized Boltzmann weight corresponding to the physical error rate is recovered at $\beta=\beta_{\mathrm{N}}$, i.e., $\pi(C)=\pi_{\beta_{\mathrm{N}}}(C)$, which follows from Eq. (11). Therefore, $\beta_{\mathrm{N}}$ provides a natural fixed-temperature choice for MCMC-based decoding [nishimori_temperature, nishimori_book, Chubb_stat_model].
@@
In Algorithm 2, to accelerate convergence of SA, we employ a logarithmic schedule for the inverse temperature, increasing $\beta$ from $\beta_{\mathrm{init}}\coloneqq 0.9\,\beta_{\mathrm{N}}$ to $\beta_{\mathrm{final}}\coloneqq\beta_{\mathrm{N}}$
@@
$\beta_{i}\coloneqq\beta_{\mathrm{init}}\,\left(1+\gamma\,\log i\right)\quad\text{for $i=1,\dots,N_{\beta}$}$ (23)
@@
$\gamma\coloneqq\begin{cases}\displaystyle\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\log N_{\beta}\right.&\text{for }N_{\beta}\geq 2,\\ 0&\text{for }N_{\beta}\in\{0,1\}.\end{cases}$ (24)
For each inverse temperature $\beta$, we execute the Metropolis step $N_{\mathrm{stabilizer}}$ times, which is counted as one Monte Carlo sweep, according to the convention.
@@
1: $S$ : measured syndrome
2: $C$ : recovery configuration, i.e., $\partial{C}=S$
@@
4:   $R\leftarrow$ greedy_matching($S$)
5:   for all $P\in\{I,X,Y,Z\}$ do
6:    $C_{L_{P}}\coloneqq R\,L_{P}$ $\triangleright$ Initial configuration for SA
@@
8:      $E^{(i)}_{L_{P}}\leftarrow$ simulated_annealing(copy($C_{L_{P}}$))
9:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}} \triangleright$ estimate of $E^{\min}_{L_{P}}$
@@
12:   $P^{\ast}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (21) 13:
14:   return $C_{L_{P^{\ast}}}$ $\triangleright$ defined as $R\,L_{P^{\ast}}$ in Line 6
@@
2: $E_{\mathrm{min}}$ : an estimate of the minimum energy encountered during annealing
@@
5:
$E_{\mathrm{min}}\leftarrow E$
6:
7:   for $i=1$ to $N_{\beta}$ do
8:    $\beta\leftarrow\beta_{\mathrm{init}}\,\left(1+\gamma\,\log{i}\right)$ $\triangleright$ Eq. (23)
9:    repeat $N_{\mathrm{stabilizer}}$ times do $\triangleright$ one Monte Carlo sweep
10:      $\Delta E\leftarrow$ Metropolis_step($\beta$,  $C$)
11:      $E\leftarrow E+\Delta E$
12:      $E_{\mathrm{min}}\leftarrow\min\{E_{\mathrm{min}},\,E\}$ $\triangleright$ Update min energy
13:
14:   return $E_{\mathrm{min}}$
Algorithm 3 Metropolis method that applies a stabilizer generator
1:Input: $\beta$ : inverse temperature
2:Input: $C$ : error configuration (modified in place)
3: Energy difference of error configurations
4: function Metropolis_step($\beta$,  $C$)
5:   Select a stabilizer generator $G_{f}$ uniformly at random.
6:   $\Delta E\leftarrow\mathcal{H}(C\,G_{f})-\mathcal{H}(C)$ 7:
8:   Draw $r\sim\mathrm{Uniform}(0,1)$.
9:   if $r<\min\{1,\exp(-\beta\,\Delta E)\}$ then
10:    $C\leftarrow C\,G_{f}$ $\triangleright$ Applies $G_{f}$ (by bit flips)
11:    return $\Delta E$
12:   else
13:    return 0